OpenStudy (anonymous):

a = 40, c = 41, b = 9 57 81

OpenStudy (anonymous):

@RockerSk8er

OpenStudy (anonymous):

Do the same thing I did last time, do you remember what it was?

OpenStudy (anonymous):

is it 81?

OpenStudy (anonymous):

That's wrong, show me your work so I can show you here you messed up. :)

OpenStudy (anonymous):

i got the answer:) thnx

OpenStudy (anonymous):

What was it?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

Again, show your work.

OpenStudy (anonymous):

b= sqrt(41^2 - 40^2) =9

OpenStudy (anonymous):

That's not the right formula. \[a^2 + b^2 = c^2\] input your a and b

OpenStudy (anonymous):

\[a^2 = 40^2\] \[b^2 = 41^2\] \[40^2 = ?\] \[41^2 = ?\]